Campbell Manufacturing Company uses two departments to make its products. Department I is a cutting department that is machine intensive and uses very few employees. Machines cut and form parts and then place the finished parts on a conveyor belt that carries them to Department II, where they are assembled into finished goods. The assembly department is labor intensive and requires many workers to assemble parts into finished goods. The company’s manufacturing facility incurs two significant overhead costs: employee fringe benefits and utility costs. The annual costs of fringe benefits are $264,000 and utility costs are $192,000. The typical consumption patterns for the two departments are as follows:

Department I Department II Total
Machine hours used 15,700 4,300 20,000
Direct labor hours used 5,300 10,700 16,000


The supervisor of each department receives a bonus based on how well the department controls costs. The company’s current policy requires using a single allocation base (machine hours or labor hours) to allocate the total overhead cost of $456,000.

Required

  1. Assume that you are the supervisor of Department I. Choose the allocation base that would minimize your department’s share of the total overhead cost. Calculate the amount of overhead that would be allocated to both departments using the base that you selected.

  2. Assume that you are the supervisor of Department II. Choose the allocation base that would minimize your department’s share of the total overhead cost. Calculate the amount of overhead that would be allocated to both departments using the base that you selected.

  3. Assume that you are the plant manager and have the authority to change the company’s overhead allocation policy. Formulate an overhead allocation policy that would be fair to the supervisors of both Department I and Department II. Compute the overhead allocations for each department using your policy.

  4. Department Allocated Cost
    I
    II
    Total
  5. Department Allocated Cost
    I
    II
    Total
  6. Costs Department I Department II
    Fringe benefits
    Utility
    Total
  7. Costs Department I Department II
    Fringe benefits
    Utility
    Total
Answer #1

Answer:

1. As per the Given Question We as the Supervisor of Dept. I should choose the allocation base of direct labour hours used, because it will minimize the Dept.'s Share of the Total overhead cost.

Allocation of total Overhead Cost

We will use DIrect labour hours as Cost Driver

Total Cost to be allocated= $456000

Direct Labour Hours Used= 16000

$456000/16000= $28.50 per Direct Labour Hour

Department I Overhead Cost = 5300*28.50=$151050.00

Department II Overhead Cost= 10700*28.50=$304950.00

2. As per the Given Question We as the Supervisor of Dept. II should choose the allocation base of Machine hours used, because it will minimize the Dept.'s Share of the Total overhead cost.

Allocation of total Overhead Cost

We will use Machine hours as Cost Driver

Total Cost to be allocated= $456000

Machine Hours Used= 20000

$456000/20000= $22.80 per Machine Hour

Department I Overhead Cost = 15700*22.80=$357960.00

Department II Overhead Cost= 4300*22.80=$98040.00

3. As per the Given Question We as the Plant Manager should choose the allocation base of both direct labour hours for Employee frindge benifits and Machine hours for and used, because it will minimize the Dept.'s Share of the Total overhead cost.

Allocation of total Overhead Cost

A. We will use Machine hours as Cost Driver for Utility Cost

Total Cost to be allocated= $192000

Machine Hours Used= 20000

$192000/20000= $9.60 per Machine Hour

B. We will use Direct Labour hours as Cost Driver for Employee Frindge Benifits  Cost

Total Cost to be allocated= $264000

Machine Hours Used= 16000

$264000/16000= $16.50 per Direct Labour Hour

Department I Overhead Cost = 15700*9.60+5300*16.50= $238170.00

Department II Overhead Cost= 4300*9.60+10700*16.50= $217830.00